# Soban Nasir Lone

### v1

@@ 3.5 GVI Posterior Approximation
We find the $q(\bm{{\theta}};\bm{\eta})$ within the $Q_{MFN}$ family that minimises the objective function. This class of approximate variational family is commonly used for BNNs. Once the variational family is decided, the optimisation reduces to finding the parameters, $\bm{\eta}=[\bm{\mu}^{q},\bm{\sigma}^{q}]$, of the variational family that minimises the loss, $\mathcal{L}(q)$ (Eq. 15). Thus the optimisation problem tries to find:
@@ 3.6 Training
To optimise the variational free energy objective function (acting as the loss function), as defined in Eq. 21, we employ the Bayes-by-Backprop algorithm [40]. This algorithm approximates the expectation within the objective function through Monte Carlo sampling. Specifically, we use $N_{q}=25$ samples for this approximation.
A key technique for efficient training is the reparameterisation trick [41]. This allows us to automatic differentiation to compute the unbiased gradients of the loss function with respect to the parameters and hence enables the use of gradient-based optimisers commonly employed in deep learning. In our implementation, we utilise the Adam optimiser [42] with its default settings. The entire training procedure is carried out in TensorFlow Probability [43].
@@
$\bm{{\theta}}=\bm{\mu}^{q}+\log(1+\exp(\bm{\sigma}^{q}))\odot\bm{\epsilon}$ (22)
where $\odot$ denotes element-wise multiplication. Combining the expected negative log-likelihood (Section 3.3.1) and the Rényi’s $\alpha$ divergence (Eq. 19), we obtain the final variational free energy objective:
$\mathcal{L}(q)=-\frac{1}{N_{q}}\sum_{c=1}^{N_{q}}\sum_{i=1}^{N_{1}}\sum_{k=1}^{N_{2}}\log\left(\mathcal{N}\left(s^{(i)}_{k};\mu_{k}\left(\bm{{\theta}}^{(c)}\right),\sigma_{k}\left(\bm{{\theta}}^{(c)}\right)\right)\right)+D_{AR}^{(\alpha)}\left[q\left(\bm{{\theta}}^{(c)};\bm{\eta}\right)||p\left(\bm{{\theta}}^{(c)}\right)\right],$ (23)
@@
where the $\frac{\partial\mathcal{L}(q)}{\partial\bm{{\theta}}}$ term is obtained through standard backpropagation. The overall training procedure is summarised in Algorithm 1.
@@
2:     Given training dataset: Arrange raw dataset, $\mathcal{D}$, according to Eq. 3.
@@
8:              Input training data to branch and trunk nets (Eq. 3).
9:              Obtain the predicted distribution (Section 3.2).
@@
11:         Compute the loss as per Eq. 23.
12:         Compute the gradients $\Delta_{\bm{\mu}^{q}}$ and $\Delta_{\bm{\sigma}^{q}}$ using Eq. 24 and Eq. 25.
@@ 4 Numerical studies
We investigate the influence of the hyperparameter $\alpha$ on posterior predictive performance while maintaining consistent DeepONet architecture and optimisation parameter settings across all experiments. Detailed architectural specifications for each problem are tabulated in Table 1. To optimise variational parameters, we employ a full-batch Adam optimiser with standard hyperparameter settings: maximum 10000 epochs, learning rate = 0.001, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. We execute ten independent optimisation runs with varying initialisations for each problem, discarding non-convergent cases identified by oscillatory loss behaviour.
The average values and standard deviations of the converged runs are reported for eleven different values of $\alpha$ in Table 2 and Table 3. We evaluate the model test performance using two complementary metrics: the normalised mean squared error (NMSE) and the negative log-likelihood (NLL). NMSE provides a straightforward measure of prediction error by averaging the squared differences between observed and predicted mean values of responses, normalised by the norm of the observed values. NLL, on the other hand, assesses how well the predicted probability distribution aligns with actual outcomes, accounting for both the prediction mean and its uncertainty.
@@ 4.1 Problem 1: Antiderivative Operator
where $\kappa$ is the radial basis function (RBF) covariance kernel with length-scale $\ell$. As outlined in Eq. 3, we generate a training set comprising $N_{1}=3000$ input functions $u(x)$ discretised at $M=100$ equidistant points within the domain (implying $\bm{{a}}^{(i)}\in\mathbb{R}^{100},i=1,\ldots,N_{1}$). For each input vector $\bm{{a}}^{(i)}$, we compute the corresponding solution at $N_{2}=20$ randomly selected points, i.e., $s^{(i)}(x_{k}),\;k=1,\ldots,N_{2}$. Notably, once the operator is learned, solutions can be evaluated at any arbitrary location within the domain. To assess generalisation performance, we employ a test set of $10000$ input functions, evaluating predicted solutions $s(x)$ for each of these test input functions at 100 equidistant locations in the domain.
Fig. 4 showcases the model’s predictive capabilities by visualising solutions for two representative test cases chosen to highlight its performance. This ensemble of points, representing the solution queried at the 100 equidistant locations, provides a visual approximation of the mean solution and the associated uncertainty. A comparative analysis with predictions from D-DeepONet and ground truth solutions for these representative examples further underscores the efficacy of using a different value of $\alpha$ than 1. For a comprehensive assessment, we present a comparison of performance using NMSE and NLL metrics averaged over all 10000 test cases for eleven different $\alpha$ values ranging from 0.25 to 3.00. Tables 2 and 3 present these results. In both metrics, $\alpha=1.25$ achieved the lowest NMSE and NLL. Notably, the NMSE is reduced by more than 50% compared to the standard KLD-VI (with $\alpha=1.00$), indicating a substantial improvement in mean predictions.
@@ 4.2 Problem 2: Gravity Pendulum under External Forcing
We next consider the nonlinear dynamics of a gravity pendulum subjected to an external force. The system is governed by the following nonlinear ODE:
@@
[…]
To construct the training dataset, we set $k=1$ and generate $N_{1}=3500$ training input functions, $u(t)$, sampled from the GRF defined in Eq. 27 at $M=100$ time points. For each training input function, we compute the corresponding reference solution, $s_{1}$, using a fourth-order Runge-Kutta integrator at $N_{2}=20$ randomly selected time points. This data is used to train the operator. A test set of 10000 input functions is employed to evaluate the predicted solution for each test input function at 100 equidistant points within the time domain $t\in(0,1]$. Fig. 5 presents the model’s predictions for two representative test inputs.
A visual comparison of the test prediction cases reveals that the deterministic D-DeepONet struggles to accurately capture the reference solution. The standard KLD-VI solution, obtained at $\alpha=1.00$, yields an improvement in the mean predictions, while the $\alpha$-VI DeepONet with $\alpha=2.00$ demonstrates superior performance over both in terms of mean prediction accuracy, as evident in Table 2. In terms of distributional fit, as measured by negative log-likelihood, the standard KLD-VI has the lowest NLL values and exhibits a better performance (Table 3). However, the NLL for $\alpha=2.00$ is only marginally higher. Considering both the metrics, $\alpha=2.00$ offers a favourable balance between accurate mean predictions and reasonable distributional fit.
@@ 4.3 Problem 3: Diffusion-Reaction System
To generate the training data, we numerically solve the PDE using a finite difference method on a $(x\times t)\equiv(100\times 100)$ grid, following a similar approach as in [4]. The training dataset comprises $N_{1}=500$ distinct source terms, $u(x)$, generated from a Gaussian random field with $\ell=0.5$ (Eq. 27), evaluated at $M=100$ equidistant points in space. For each training input function, the solution $s(\bm{{y}})$ is obtained at $N_{2}=100$ random points $\bm{{y}}$ sampled from the $100\times 100$ grid; these sampled solution points are denoted by filled white circles in Fig. 6. Subsequently, we employ this training dataset to learn the operator mapping from $u(x)$ to the PDE solution $s(x,t)$.
@@
To assess generalisation performance, we employ a separate test set comprising 10000 diverse source terms. Fig. 7 presents a visual comparison of the model’s predicted solution against the reference solution for a representative test source term across the entire $100\times 100$ spatiotemporal grid. The figure also includes the associated uncertainty (one standard deviation) and absolute error. It is seen that regions of low response values (visualised as bluish areas) exhibit reduced smoothness in the predicted mean solution compared to the reference. As such, these regions demonstrate higher uncertainty, as evidenced by increased standard deviation.
@@ 4.4 Problem 4: Advection-Diffusion Equation
To construct the training dataset, we generate $N_{1}=1000$ unique initial conditions by sampling $u$ from a GRF (with input domain defined by $\sin^{2}(2\pi x)$) with length-scale $\ell=0.5$ (Eq. 27), discretised at $M=100$ spatial points. The PDE is then solved numerically on a spatiotemporal grid $(x\times t)\equiv(100\times 100)$ using a finite difference method. For each initial condition, the solution $s(x,t)$ is evaluated at $N_{2}=100$ randomly sampled points within the grid, following similarly as the data generation process outlined in Fig. 6. This dataset serves to train the operator.
For evaluation, a separate test set of 10000 initial conditions is employed. Fig. 8 presents a comparison of the model’s predicted solution against the reference solution for a sample test initial condition across the entire spatiotemporal domain. A quantitative evaluation of the model’s predictive performance across different values of $\alpha$ was conducted using the NMSE and NLL metrics, summarised in Tables 2 and 3. In contrast to the previous examples where larger $\alpha$ values (i.e., values greater than 1) typically yielded superior performance, this problem exhibits optimal results at $\alpha=0.5$ in terms of both NMSE and NLL. While values of $\alpha$ greater than 1 typically led to subpar performance in this case, the specific value of $\alpha=1.75$ can be considered as a reasonable choice for $\alpha$ values greater than 1.
@@
The first OOD dataset was constructed by altering the kernel length-scale of the GRF (Eq. 27) used to generate initial conditions. While the training data utilised a length-scale of $\ell=0.5$, the OOD dataset employed a reduced length-scale of $\ell=0.2$ within the RBF kernel. This modification introduces increased fluctuations in the generated initial conditions compared to the training distribution.
@@
The predictive performance of models trained with three different $\alpha$ values: 0.5, 1.0 (standard KLD-VI), and 1.75, was evaluated on these OOD test datasets. Similar to the in-distribution analysis, the NMSE and NLL metrics were computed. The average results, summarised in Table 4, indicate that the models exhibit reasonable generalisation capabilities on OOD data. Consistent with the in-distribution findings, $\alpha=0.5$ outperformed both the standard KLD-VI ($\alpha=1.00$) and the $\alpha=1.75$ model in terms of both NMSE and NLL for both OOD datasets. This improvement over the standard KLD-VI model translated to upto a 7% reduction in NMSE and an 8.7% reduction in NLL for the tested OOD scenarios. Interestingly, the performance on the rational quadratic kernel samples was superior to that on the RBF kernel samples.
@@
RBF with $\ell=0.2$ Rational quadratic NMSE NLL NMSE NLL $\alpha$ $10^{-3}\pm 10^{-3}$ $10^{0}\pm 10^{0}$ $10^{-3}\pm 10^{-3}$ $10^{0}\pm 10^{0}$ 0.50 2.062 ± 0.269 -3.856 ± 0.069 0.197 ± 0.032 -4.352 ± 0.068 1.00 (VI) 2.207 ± 0.640 -3.547 ± 0.414 0.212 ± 0.040 -4.093 ± 0.358 1.75 2.585 ± 0.807 -3.752 ± 0.095 0.210 ± 0.056 -4.266 ± 0.154

### v2

@@ 3.5 GVI posterior approximation
We find the $q(\bm{{\theta}};\bm{\eta})$ within the $Q_{MFN}$ family that minimises the objective function. This class of approximate variational family is commonly used for BNNs and is the same as that used in VB-DeepONet [garg2023vb]. Once the variational family is decided, the optimisation reduces to finding the parameters, $\bm{\eta}=[\bm{\mu}^{q},\bm{\sigma}^{q}]$, of the variational family that minimises the loss, $\mathcal{L}(q)$ (Eq.˜15). Thus the optimisation problem tries to find:
@@ 3.6 Training
To optimise the variational free energy objective function (acting as the loss function), as defined in Eq.˜21, we employ the Bayes-by-Backprop algorithm [blundell2015weight]. This algorithm approximates the expectation within the objective function through Monte Carlo sampling. Specifically, we use $N_{q}=25$ samples for this approximation.
A key technique for efficient training is the reparameterisation trick [kingma2015variational]. This allows us to use automatic differentiation to compute the unbiased gradients of the loss function with respect to the parameters and hence enables the use of gradient-based optimisers commonly employed in deep learning. In our implementation, we utilise the Adam optimiser [kingma2014adam] with its default settings. The entire training procedure is carried out in TensorFlow Probability [dillon2017tensorflow].
@@
$\bm{{\theta}}=\bm{\mu}^{q}+\log(1+\exp(\bm{\sigma}^{q}))\odot\bm{\epsilon},$ (22)
where $\odot$ denotes element-wise multiplication. Combining the expected negative log-likelihood (Section˜3.3.1) and the Rényi’s $\alpha$-divergence (Eq.˜19), we obtain the final variational free energy objective:
$\mathcal{L}(q)=-\frac{1}{N_{q}}\sum_{c=1}^{N_{q}}\sum_{i=1}^{N_{1}}\sum_{k=1}^{N_{2}}\log\left(\mathcal{N}\left(s^{(i)}_{k};\mu_{k}\left(\bm{{\theta}}^{(c)}\right),\sigma_{k}\left(\bm{{\theta}}^{(c)}\right)\right)\right)+D_{AR}^{(\alpha)}\left[q\left(\bm{{\theta}}^{(c)};\bm{\eta}\right)\mid\mid p\left(\bm{{\theta}}^{(c)}\right)\right],$ (23)
@@
where the $\frac{\partial\mathcal{L}(q)}{\partial\bm{{\theta}}}$ term is obtained through standard backpropagation. The overall training procedure is summarised in Algorithm˜1.
@@
2:  Given training dataset: Arrange raw dataset, $\mathcal{D}$, according to Eq.˜3.
@@
8:     Input training data to branch and trunk nets (Eq.˜3).
9:     Obtain the predicted distribution (Section˜3.2).
@@
11:   Compute the loss as per Eq.˜23.
12:   Compute the gradients $\Delta_{\bm{\mu}^{q}}$ and $\Delta_{\bm{\sigma}^{q}}$ using Eq.˜24 and Eq.˜25.
@@ 4 Numerical studies
We investigate the influence of the hyperparameter $\alpha$ on posterior predictive performance while maintaining consistent DeepONet architecture and optimisation settings across all experiments. Detailed architectural specifications for each problem are tabulated in Table˜1.
Our experimental protocol employs full-batch Adam optimisation with fixed hyperparameters across all methods: learning rate = 0.001, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, and a maximum of 10,000 epochs. This simplified setup isolates the effect of $\alpha$ on the quality of inference. We note that standard techniques – such as mini-batch training, learning rate scheduling, and explicit regularisation – can substantially improve all DeepONet variants [lu2021learning, lu2022comprehensive], but are not employed here to ensure a fair comparison focused solely on the divergence measure. We execute ten independent optimisation runs with varying initialisations for each problem, discarding non-convergent cases identified by oscillatory loss behaviour.
The average values and standard deviations of the converged runs are reported for eleven different values of $\alpha$ in Table˜2 and Table˜3. We evaluate the model test performance using two complementary metrics: the normalised mean squared error (NMSE) and the negative log-likelihood (NLL). NMSE provides a normalised measure of prediction error by computing the average squared difference between observed values and the predicted mean values of responses. It is then scaled by the squared mean of the observed values to ensure a relative, unit-independent error measure. This allows for a meaningful comparison of prediction accuracy across different datasets and scales. NLL, on the other hand, assesses how well the predicted probability distribution aligns with actual outcomes, accounting for both the prediction mean and its uncertainty.
@@ 4.1 Problem 1: Antiderivative operator
where $\kappa$ is the radial basis function (RBF) covariance kernel with length-scale $\ell$. As outlined in Eq.˜3, we generate a training set comprising $N_{1}=3000$ input functions $u(x)$, with $\ell=0.5$, discretised at $M=100$ equidistant points within the domain (implying $\bm{{a}}^{(i)}\in\mathbb{R}^{100},i=1,\ldots,N_{1}$). For each input vector $\bm{{a}}^{(i)}$, we compute the corresponding solution at $N_{2}=20$ randomly selected points, i.e., $s^{(i)}(x_{k}),\;k=1,\ldots,N_{2}$. Notably, once the operator is learned, solutions can be evaluated at any arbitrary location within the domain. To assess generalisation performance, we employ a test set of $10,000$ input functions, evaluating predicted solutions $s(x)$ for each of these test input functions at 100 equidistant locations in the domain.
Fig.˜4 showcases the model’s predictive capabilities by visualising solutions for two representative test cases chosen to highlight its performance. This ensemble of points, representing the solution queried at the 100 equidistant locations, provides a visual approximation of the mean solution and the associated uncertainty. A comparative analysis with predictions from D-DeepONet and ground truth solutions for these representative examples further underscores the efficacy of using a different value of $\alpha$ than 1. For a comprehensive assessment, we present a comparison of performance using NMSE and NLL metrics averaged over all 10,000 test cases for eleven different $\alpha$ values ranging from 0.25 to 3.00. Tables 2 and 3 present these results. In both metrics, $\alpha=1.25$ achieved the lowest NMSE and NLL. Notably, the NMSE is reduced by more than 50% compared to the standard KLD-VI (with $\alpha=1.00$), indicating a substantial improvement in mean predictions.
@@ 4.2 Problem 2: Gravity pendulum under external forcing
Next, we consider the non-linear dynamics of a gravity pendulum subjected to an external force. The system is governed by the following non-linear ODE:
@@
with initial conditions $\bm{{s}}(0)=\left[0\;\;0\right]^{T}$, where $s_{1}=\phi$ and $s_{2}=\frac{\mathrm{d}\phi}{\mathrm{d}t}$. Given the purely time-dependent nature of the problem. the domain reduces to a scalar temporal variable, i.e., $y=t$. For this problem, we consider the generalised input function $a(y)$ as equal to the time-dependent forcing function $u(t)$.
To construct the training dataset, we set $k=1$ and generate $N_{1}=3500$ training input functions, $u(t)$, sampled from the GRF defined in Eq.˜27 at $M=100$ time points. For each training input function, we compute the corresponding reference solution, $s_{1}$, using a fourth-order Runge-Kutta integrator at $N_{2}=20$ randomly selected time points. This data is used to train the operator. A test set of 10,000 input functions is employed to evaluate the predicted solution for each test input function at 100 equidistant points within the time domain $t\in(0,1]$. Fig.˜5 presents the model’s predictions for two representative test inputs.
A visual comparison of the test prediction cases reveals that the deterministic D-DeepONet struggles to accurately capture the reference solution. The standard KLD-VI solution, obtained at $\alpha=1.00$, yields an improvement in the mean predictions, while the $\alpha$-VI DeepONet with $\alpha=2.00$ demonstrates superior performance over both in terms of mean prediction accuracy, as evident in Table˜2. In terms of distributional fit, as measured by negative log-likelihood, the standard KLD-VI has the lowest NLL values and exhibits a better performance (Table˜3). However, the NLL for $\alpha=2.00$ is only marginally higher. Considering both the metrics, $\alpha=2.00$ offers a favourable balance between accurate mean predictions and reasonable distributional fit.
@@ 4.3 Problem 3: Diffusion-reaction system
To generate the training data, we numerically solve the PDE using a finite difference method on a $(x\times t)\equiv(100\times 100)$ grid, following a similar approach as in [lu2021learning]. The training dataset comprises $N_{1}=500$ distinct source terms, $u(x)$, generated from a Gaussian random field with $\ell=0.5$ (Eq.˜27), evaluated at $M=100$ equidistant points in space. For each training input function, the solution $s(\bm{{y}})$ is obtained at $N_{2}=100$ random points $\bm{{y}}$ sampled from the $100\times 100$ grid; these sampled solution points are denoted by filled white circles in Fig.˜6. Subsequently, we employ this training dataset to learn the operator mapping from $u(x)$ to the PDE solution $s(x,t)$.
@@
To assess generalisation performance, we employ a separate test set comprising 10,000 diverse source terms. Fig.˜7 presents a visual comparison of the model’s predicted solution against the reference solution for a representative test source term across the entire $100\times 100$ spatiotemporal grid. The figure also includes the associated uncertainty (one standard deviation) and absolute error. It is seen that regions of low response values (visualised as bluish areas) exhibit reduced smoothness in the predicted mean solution compared to the reference. As such, these regions demonstrate higher uncertainty, as evidenced by increased standard deviation.
@@ 4.4 Problem 4: Advection-diffusion equation
To construct the training dataset, we generate $N_{1}=1000$ unique initial conditions by sampling $u$ from a GRF (with input domain defined by $\sin^{2}(2\pi x)$) with length-scale $\ell=0.5$ (Eq.˜27), discretised at $M=100$ spatial points. The PDE is then solved numerically on a spatiotemporal grid $(x\times t)\equiv(100\times 100)$ using a finite difference method. For each initial condition, the solution $s(x,t)$ is evaluated at $N_{2}=100$ randomly sampled points within the grid, following similarly as the data generation process outlined in Fig.˜6. This dataset serves to train the operator.
For evaluation, a separate test set of 10,000 initial conditions is employed. Fig.˜8 presents a comparison of the model’s predicted solution against the reference solution for a sample test initial condition across the entire spatiotemporal domain. A quantitative evaluation of the model’s predictive performance across different values of $\alpha$ was conducted using the NMSE and NLL metrics, summarised in Tables 2 and 3. In contrast to the previous examples where larger $\alpha$ values (i.e., values greater than 1) typically yielded superior performance, this problem exhibits optimal results at $\alpha=0.5$ in terms of both NMSE and NLL. While values of $\alpha$ greater than 1 typically led to subpar performance in this case, the specific value of $\alpha=1.75$ can be considered as a reasonable choice for $\alpha$ values greater than 1.
@@ 4.5 Out-of-distribution generalisation
The first OOD dataset was constructed by altering the kernel length-scale of the GRF (Eq.˜27) used to generate initial conditions. While the training data utilised a length-scale of $\ell=0.5$, the OOD dataset employed a reduced length-scale of $\ell=0.2$ within the RBF kernel. This modification introduces increased fluctuations in the generated initial conditions compared to the training distribution.
@@
The predictive performance of models trained with three different $\alpha$ values: 0.5, 1.0 (standard KLD-VI), and 1.75, was evaluated on these OOD test datasets. Similar to the in-distribution analysis, the NMSE and NLL metrics were computed. The average results, summarised in Table˜4, indicate that the models exhibit reasonable generalisation capabilities on OOD data. Consistent with the in-distribution findings, $\alpha=0.5$ outperformed both the standard KLD-VI ($\alpha=1.00$) and the $\alpha=1.75$ model in terms of both NMSE and NLL for both OOD datasets. This improvement over the standard KLD-VI model translated to up to a 7% reduction in NMSE and an 8.7% reduction in NLL for the tested OOD scenarios. Interestingly, the performance on the rational quadratic kernel samples was superior to that on the RBF kernel samples.
@@
RBF with $\ell=0.2$ Rational quadratic NMSE NLL NMSE NLL $\alpha$ $^{-3}\pm 10^{-3} 100±1$$^{0}\pm 10^{0}$ $^{-3}\pm 10^{-3} 100±1$$^{0}\pm 10^{0}$ 0.50 2.062 ± 0.269 -3.856 ± 0.069 0.197 ± 0.032 -4.352 ± 0.068 1.00 (VI) 2.207 ± 0.640 -3.547 ± 0.414 0.212 ± 0.040 -4.093 ± 0.358 1.75 2.585 ± 0.807 -3.752 ± 0.095 0.210 ± 0.056 -4.266 ± 0.154
4.6 Robustness to noisy observations
To assess the robustness of the proposed $\alpha$–VI DeepONet under more realistic conditions, we conducted additional experiments by contaminating the training data with Gaussian noise of controlled magnitude. This setup emulates practical scenarios where sensor readings or experimental measurements are subject to random perturbations. For each problem, the clean training data $s^{(i)}(y_{k})$ were perturbed as:
$s_{\text{noisy}}=s+\epsilon,\qquad\epsilon\sim\mathcal{N}(0,\sigma_{\text{noise}}^{2}),$ (29)
where the noise standard deviation was set relative to the clean signal’s standard deviation as
$\sigma_{\text{noise}}=\frac{p}{100}\cdot\operatorname{std}(s),$ (30)
with $p$ the noise level in percent. We tested two representative problems:
  • Gravity pendulum under external forcing, with $p\in\{5\%,\,10\%\}$
  • Diffusion-reaction system, with $p=10\%$
Three approaches are compared: (i) D-DeepONet, (ii) variational DeepONet with KL divergence (KLD–VI, $\alpha=1.00$), and (iii) the proposed $\alpha$–VI DeepONet with the best-performing $\alpha$ identified on the training set ($\alpha=2.00$ and $\alpha=3.00$ respectively). Tables 5 and 6 summarise the NMSE and NLL on the corresponding test sets.
The proposed $\alpha$-VI DeepONet demonstrates superior noise robustness compared to standard variational inference, maintaining competitive performance relative to noise-free conditions across both NMSE and NLL metrics. In the gravity pendulum problem, the best noise-free NMSE performer ($\alpha=2.00$, NMSE = 1.111 $\times 10^{-4}$) shows minimal degradation under 5% noise (NMSE = 1.190 $\times 10^{-4}$), while outperforming standard VI (NMSE = 1.370 $\times 10^{-4}$) under identical conditions. The NLL results corroborate this trend, with optimal configurations maintaining superior likelihood performance compared to standard VI across noise conditions. Under 10% noise, while standard VI performs better than $\alpha=2.00$ in NMSE, both methods demonstrate reasonable performance relative to their noise-free baselines. In the diffusion-reaction problem, the optimal noise-free configuration ($\alpha=3.00$, NMSE = 4.463 $\times 10^{-3}$, NLL = 3.658-3.658) exhibits strong resilience under 10% noise (NMSE = 5.867 $\times 10^{-3}$, NLL = 3.575-3.575), significantly outperforming standard VI ($\alpha=1.00$: NMSE = 7.618 $\times 10^{-3}$, NLL = 3.439-3.439) across both metrics. These consistent patterns across NMSE and NLL demonstrate that the $\alpha$-VI framework provides enhanced flexibility in handling observational uncertainty, maintaining robust probabilistic predictions suitable for real-world applications where measurement uncertainties are inevitable.
Table 5: Performance under noisy observations for the gravity pendulum under external forcing problem. The best-performing model for each metric is highlighted in bold.
5% Noise 10% Noise NMSE NLL NMSE NLL $\alpha$ $^{-4}\pm 10^{-4} 100±1$$^{0}\pm 10^{0}$ $^{-4}\pm 10^{-4} 100±1$$^{0}\pm 10^{0}$ D–DeepONet 2.609 $\pm$ 0.920 3.410 $\pm$ 1.680 1.00 (VI) 1.370 $\pm$ 0.370 $4.527\pm 0.005$ $\bm{2.090\pm 0.270}$ $\bm{-3.943\pm 0.003}$ 2.00 $\bm{1.190\pm 0.180}$ $\bm{-4.536\pm 0.010}$ 2.450 $\pm$ 0.080 $3.942\pm 0.005$
Table 6: Performance under noisy observations for the diffusion-reaction problem. The best-performing model for each metric is highlighted in bold.
10% Noise NMSE NLL $\alpha$ $^{-3}\pm 10^{-3} 100±1$$^{0}\pm 10^{0}$ D-DeepONet 8.620 $\pm$ 1.710 1.00 (VI) 7.618 $\pm$ 2.154 $3.439\pm 0.134$ 3.00 $\bm{5.867\pm 0.609}$ $\bm{-3.575\pm 0.035}$